OpenStudy (anonymous):

for the power series (2x)^n/n^4 find the radius and integral of convergence.

OpenStudy (anonymous):

this is what i've done and am getting the wrong answer...

OpenStudy (anonymous):

\[\frac{ 2x ^{n+1} }{ (n+1)^{4} }*\frac{ n ^{4} }{ (2x)^{n} }\] \[\lim_{n \rightarrow \inf}\left| \frac{ 2xn ^{4} }{ (n+1)^{4} } \right|\] \[\left| 2x \right|\lim_{n \rightarrow \inf}\frac{ n ^{4} }{ (n+1)^{4} }\] \[\left| 2x \right|\] so \[-1<2x <1\] meaning R=1 solve for x... integral is -1/2 to 1/2. but Webwork says that's wrong.....

OpenStudy (lukecrayonz):

@Directrix do you know how to do this? >_<

OpenStudy (lukecrayonz):

Very funny, but I don't know how which is why I'm asking you haha

OpenStudy (lukecrayonz):

She's been on for an hour just waiting which is why I was asking, I never learned this and was googling things to learn it but I still don't know.

OpenStudy (paounn):

Careful reading brackets, second denominator is 2^n * x^n

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

i wrote the top part wrong, but did it right. it's supposed to be \[\frac{ (2x)^{(n+1)} }{ (n+1)^{4} }*\frac{ n ^{4} }{ (2x)^{n} }\]

OpenStudy (anonymous):

so the n+1-n still cancel each other leaving 2x at the top.

OpenStudy (paounn):

then your radius should be\[\lim_{x \rightarrow \infty} \left| a_n \over a_{n+1} \right|\] if it exists.

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/Classes/CalcII/PowerSeries.aspx this says the radius is \[\frac{ a_{n+1} }{ a _{n} }\]

OpenStudy (anonymous):

so i'm confused. which is it?

OpenStudy (paounn):

Ups, my mistake, got a bit rusty, it's yours. :)

OpenStudy (amistre64):

(n+1)/n or n/(n-1) :)

OpenStudy (anonymous):

ok, so since that's what i did, why is webwork saying i got the wrong answer?

OpenStudy (amistre64):

how do you determine that your results are wrong?

OpenStudy (amistre64):

can you take a screen shot ?

OpenStudy (anonymous):

yeah, hang on

OpenStudy (amistre64):

try a radius of .5, since you have unlimited attempts :)

OpenStudy (anonymous):

that was correct....so why was my radius .5 and not 1?

OpenStudy (amistre64):

\[|2x|<1\] \[2|x|<1\] \[|x|<1/2\]

OpenStudy (paounn):

Think the way a "circle" radius work, it goes from the center to the side, not that one side to the opposite :)

OpenStudy (anonymous):

so the radius is gonna be the absolute value of the endpoint?

OpenStudy (amistre64):

if your interval ... the diameter, is from -.5 to .5; the radius (half the diameter) is .5 in length

OpenStudy (amistre64):

in general; \[|f(x-a)|~\lim~c\] \[R=\frac1c\]

OpenStudy (amistre64):

you should only pull out the naked x parts, and leave everything else inside the limit. So pulling out the 2 alongside the x was a bit off putting i think

OpenStudy (amistre64):

and as Paounn stated, you cant really have a radius that is larger than half the interval

OpenStudy (anonymous):

ok. i think i understand. maybe. we'll see when i attempt the next problem...thanks amistre64 and paounn. you've been very helpful.

OpenStudy (amistre64):

good luck ;)